OpenStudy (anonymous):

you have 10 fewer quarters than dimes and 5 fewer nickels than quarters. the total value of the coins is 4.75. how many quarters, dimes and nickels do you have? please tell me how you set up the equations

OpenStudy (anonymous):

First I would start with 0.05n + 0.10d + 0.25q = 4.75 Are you okay with this?

OpenStudy (anonymous):

sure but i need 3 equations to solve for all three variables

OpenStudy (anonymous):

From the first sentence, second equation is 10 + q = d Can you get the third equation from the remainder of the first sentence?

OpenStudy (anonymous):

0.05[(d-10)-5] + 0.10(d) + 0.25(d-10) = 4.75 Then solve for d

OpenStudy (anonymous):

5+n=q?

OpenStudy (anonymous):

Yes! Need further help?

OpenStudy (anonymous):

ill let you know

OpenStudy (anonymous):

i got 5 nickels, is that right?

OpenStudy (anonymous):

Let me check...

OpenStudy (anonymous):

20 dimes and 10 quarters?

OpenStudy (anonymous):

0.05(5) + 0.10(20) + 0.25(10) = 0.25 + 2 + 2.50 = 4.75 Nice job!!

OpenStudy (anonymous):

ahhh thank youu!!